how many nitrogen molecules are in a box measuring 9.00cm H x 21.00cm L x 14.50cm W


1
Expert's answer
2022-02-28T11:01:27-0500

Volume= 9×21×14.5= 2740.5cm3

Mass=2740.5g

Moles= 2740.5/14= 195.75

Avagadro constant= 6.022× 10^23 mol-1

Molecules= 6.022× 10^23 × 195.75= 1.18× 10^26